Reports 1-1 of 1 Clear search Modify search
takafumi.ushiba - 23:14 Thursday 02 February 2023 (23815) Print this report
ADS of OMMT2 and OSTM for fine alignment to OMC

[Akutsu (remote), Ushiba]

We engaged ADSs to OMMT2 and OSTM for fine alignment to OMC.
It seems working well and transmitted power is increased by a factor of 1.3.

Detail will be posted tomorrow.

During this work, we reduced laser power as already repoted in klog23814.
During the work, we modified OMC_ASC and OMC_LSC guardian slightly.
It doesn't affect the OMC lock but if you find any porblems, please let us know.

Comments to this report:
tomotada.akutsu - 11:21 Friday 03 February 2023 (23818) Print this report

In the middle of the work, the OMC_ASC guardian threw an error on python-threading and die, and we spent a great deal of time working around it; not resolved yet but we just somehow "omit" it now. It appears to be due to a conflict among an already-existing thread and newly started one.

Although this happened today largely due to an unfortunate coincidence or other factors (like some unexpected lines in the OMC_LSC guardian, so we modified them then), nevertheless, it means that there is an error mixed in with the paths that could be taken, so one should do modify it.

takafumi.ushiba - 12:31 Friday 03 February 2023 (23820) Print this report


First, I implemented OMMT2T centering loop to OMMT1, which is reported in klog23728.
Then, We tested ADSs of OMMT2 and OSTM for fine alignment to OMC.
Since the OMC QPD loop is currently stable (it was not stable during O3 due to glitch problems), we feedback ADS signals to the setpoint of QPD loops.

Followings are the detail of setting of ADSs:
Dither frequency: 15 Hz, 16 Hz, 17 Hz, and 18 Hz for OMMT2P, OMMT2Y, OSTMY, and OSTMP, respectively.
Dither amplitude: 200 cnts for pitch and 600 cnts for yaw.
INF filter: FM1 (BandPass filter)
OUTF filter: FM1 (LP1), FM2 (comb15), FM3 (comb1), FM9 (integrator)

Figure 1 shows the time series of ADS feedback and OMC output.
Transmission power increases from 34 to 50, and ADSs seem working well.
I implemnted the ADSs in INITIAL_ALIGNMENT guardian but maybe it is better to move it some other guardian like OMC_ASC.


During the work, OMC_ASC guardian moved OMMT1 and kicked the suspension because it was used for OMMT2T QPD centering.
Since OMMT1 doesn't use OMC alignment, I removed the request to OMMT1 from OMC_ASC guardian not to conflict the alignment control.

Also, we faced some bugs of OMC_ASC guardian.
Figure 2 shows the log around the bug.
It seems new thread is called withouut killing the existing thread, so it needs to be fixed.

Images attached to this comment
tomotada.akutsu - 12:43 Friday 03 February 2023 (23821) Print this report

By the way, is this transmission power of OMC consistent? Maybe yes as follows, but please someone check it precisely:

  • IMC output around 2.2 W
  • PRM reduces it 1/10, so it becomes 220 mW.
  • Before reaching OMC, the beam passes BS twice, so 1/2*1/2 = 1/4, meaning 55 mW.
takafumi.ushiba - 21:49 Sunday 05 February 2023 (23842) Print this report

I moved OMMT2T QPD centering loop to ASC_LOCK guardian and OMC ADS to OMC_ASC guardian.
I also modified the initial alignment guardian and test it.
They are working well, so I will prepare the document soon.

Since the number of requestable state is over 15 in INITIAL_ALIGNMENT guardian, we cannot request new state from the nominal guardian medm.
we can request them from the "all" button but it is better to prepare the special medm like the other initial alignment procedure.

Search Help